jowhite:

An elephant’s heart rate was measured at 26 beats in 1 minute. If the heart rate remains the same, how many times would the elephant’s heart beat in 4 minutes? A.104 beats B.78 beats C.130 beats

ghostly111:

The answer A. 104 beats

ghostly111:

So 26 beats in one minute and then you asked how many beats in 4 minutes I multiplied 26 times 4 and I got 104

montes67:

well i mean if 1=26 that would be like 4 *26= which is....4*6 = 24 and then 2*4 = 8 +2 =10 sooo that would be 104. 4 min= 104 heartbeats